# Autoregressive models for panel data causal inference with application to state-level opioid policies

### v1

@@ 3.1 Additional time dependence
The main extensions we present allow for greater dependence of the variables over time, which allows us to relax the no carryover effects and ignorability assumptions. First consider no carryover effects: in many policy settings, treatment in earlier time periods likely affects outcomes at time $t$, and therefore, the effects of entire treatment sequences are of interest. To incorporate these effects we first redefine our causal estimand, now given by $\Delta_{t}(\bar{a})=\mathbb{E}[Y_{it}(\bar{a})-Y_{it}(\bar{0})].$ This parameter represents the effect of a particular treatment sequence compared against no treatment at time $t$. While in principle this allows for effects of all prior treatments, in practice, we make the assumption that this is limited to only the most recent $l$ time points.
@@
Let $\bar{a}$ and $\bar{a}^{\prime}$ be two treatment sequences such that $a_{t}=a_{t}^{\prime}$, $a_{t-1}=a_{t-1}^{\prime},\dots,a_{t-l}=a_{t-l}^{\prime}$, then we have $Y_{it}(\bar{a})=Y_{it}(\bar{a}^{\prime}),\ \forall\bar{a}.$
@@
Second, we relax Assumption 3 to condition on $k$ prior time period’s counterfactuals as opposed to only the single time period’s counterfactual outcome used previously. This relaxation increases the plausibility of the assumption, particularly in settings where we worry about the dependence of the outcomes over time. However, in order to relax this assumption, we must also strengthen the positivity assumption (Assumption 5) to include these additional time-periods. We can then incorporate these relaxations into our estimation strategy by fitting the following version of the L-DAM:
$Y_{it}$ $=\alpha+\sum_{b=1}^{k}\left\{\delta_{b}\left(Y_{i,t-b}-\sum_{z=0}^{l}\theta_{z}A_{i,t-b-z}\right)\right\}+\sum_{z=0}^{l}\theta_{z}A_{i,t-z}+\epsilon_{it}$
It is easy to see how this model incorporates these extensions. First, lagged policy effects are now given by the $\theta_{z}$ parameters. Additionally, the model includes linear adjustments for estimates of $Y_{i,t-1}(0),\dots,Y_{i,t-k}(0)$ instead of estimates of just $Y_{i,t-1}(0)$.
@@ 3.2 Time-varying covariates
We can also relax our ignorability assumption to allow for both time-invariant and time-varying covariates. Assume, for example, that covariates $X_{it}$ are determined prior to $Y_{it}$. Then we could assume that ignorability holds conditional on covariates and $k$ lagged potential outcomes absent treatment $(X_{it}$, $Y_{it-k}(0),...,Y_{it-1}(0))$. However, we must again change the positivity requirement to further condition on these quantities. One can then incorporate these same covariates into the L-DAM as:
$Y_{it}$ $=\alpha+\sum_{b=1}^{k}\left\{\delta_{b}\left(Y_{i,t-b}-\sum_{z=0}^{l}\theta_{z}A_{i,t-b-z}\right)\right\}+\sum_{z=0}^{l}\theta_{z}A_{i,t-z}+\Gamma^{\top}X_{it}+\epsilon_{it}$
When the covariates $X_{it}$ are time-varying, implicit in this specification are that the covariates are exogenous, or unaffected by treatment. That is, we further assume,
@@ 3.3 Multiple policies
A separate but important type of time-varying variable we may also observe are other policy indicators $M_{it}$. For example, state legislatures may consider multiple policies to tackle opioid-related outcomes, including both naloxone access laws and Medicaid prescription coverage policies. In this situation, we may wish to know the effect of one policy, holding the other policy fixed at some value. In the language of potential outcomes, we are interested in a so-called “controlled direct effect” (Nguyen, Schmid and Stuart, 2021).To be precise, assume that in each time period states may assign either policies $A$ or $M$, but we are primarily interested in the effects of policy $A$. One relevant causal quantity would then be $\psi_{t}(\bar{a},\bar{m})=\mathbb{E}[Y_{it}(\bar{a},\bar{m})-Y_{it}(\bar{0},\bar{m})]$, where potential outcomes are now indexed by both treatment values. Identifying this quantity in terms of the observed data distribution requires analogues to Assumptions (1)-(5). We formally describe these assumptions in Appendix A; however, intuitively we can view $(A,M)$ as a multi-valued treatment randomly assigned conditional on the prior potential outcomes absent any policy. [^3] We can allow that the randomization for $M$ is conditional $A$ or vice versa. In much of the mediation literature it is assumed that $A$ is determined prior to $M$. We instead make no assumption on their causal ordering, at the price of a slightly stronger conditional independence assumption. Under these assumptions we can identify the causal estimand in terms of the observed data distribution as an average of a conditional expectation function, described in Appendix A. For continuous outcomes we can again parameterize a model for this function using an L-DAM:
$Y_{it}$ $=\alpha+\sum_{b=1}^{k}\left\{\delta_{b}\left(Y_{i,t-b}-\sum_{z=0}^{l}[\theta_{z}A_{i,t-b-z}+\zeta_{z}M_{i,t-b-z}]\right)\right\}+\sum_{z=0}^{l}[\theta_{z}A_{i,t-z}+\zeta_{z}M_{i,t-z}]+\epsilon_{it}$
@@ 3.4 Nonlinear extensions
Linear models may be inappropriate for modeling the conditional expectation function for certain outcomes. For example, in opioid studies, we are often faced with count data, where a log-link function may be more appropriate. We can continue to invoke Assumptions (1)-(5) above, but in place of assumption (4), we may instead assume constant multiplicative treatment effects:
@@
Analogous to the L-DAM, we can then parameterize the conditional expectation function with respect to a log-link and assuming $\eta_{t}(\bar{a};\theta_{t})=\theta^{\top}\bar{a}$:
$\log(\mathbb{E}[Y_{it}\mid\bar{Y}_{i,t-1},\bar{A}_{it}])$ $=\alpha+\sum_{b=1}^{k}\left\{\delta_{b}\left(\log(Y_{i,t-b})-\sum_{z=0}^{l}\theta_{z}A_{i,t-1-z}\right)\right\}+\sum_{z=0}^{l}\theta_{z}A_{i,t-z}.$
@@
$\tilde{\Delta}_{t}(\bar{a})=\frac{\mathbb{E}[Y_{it}(\bar{a})]}{\mathbb{E}[Y_{it}(\bar{0})]}.$
@@ 4 Connections to existing identification strategies
In this section we briefly compare our identification strategy to other strategies commonly used in causal inference with panel data. We first compare our approach to difference-in-differences and synthetic controls, commonly used in applied policy applications, and we then compare it to the sequential g-formula, which identifies causal quantities under sequentially randomized trials and is a more commonly used approach in epidemiologic studies.
@@ 4.1 Difference-in-differences and synthetic controls
As mentioned before, difference-in-differences and synthetic controls conventionally target the treatment effect on the treated (ATT), while our identification strategy can be used to target the ATT or the average treatment effect (ATE). Because the potential outcomes under treatment are fully observed, for DiD and synthetic controls analyses, the only quantity requiring causal identification is the mean potential outcomes absent treatment among the treated group. Synthetic controls do this by assuming ignorability of the potential outcomes absent treatment conditional on pre-treatment outcomes. [^4] Alternatively, synthetic controls have frequently been motivated by assuming that the potential outcomes absent treatment follow a linear factor model (Abadie, Diamond and Hainmueller, 2010). By contrast, DiD assumes “parallel-trends:” that the mean trends in the untreated group equal the counterfactual trends absent treatment in the treated group. This assumption allows for selection into treatment to be based on an unmeasured variable causing a time-invariant difference between the mean outcomes in the treatment and control groups.
To compare these assumptions to ours, consider the canonical DiD setting with two time-periods ($t=1,2$), where treatment only occurs in the second time-period. In this case, our ignorability assumption simply states that treatment is randomized with respect to the potential outcomes in the second time-period conditional on the first-period outcomes. Moreover, this assumption implies the ignorability assumption justifying synthetic controls Abadie, Diamond and Hainmueller (2010); Ben-Michael, Feller and Rothstein (2021). One key difference, is that our assumption is stronger: we require ignorability for both potential outcomes – $Y_{i2}(1),Y_{i2}(0)$ – while synthetic controls only requires an ignorability assumption with respect to $Y_{i2}(0)$. Beyond this, any comparison between conditional randomization and parallel trends in this setting immediately applies to our approach. In a more general setting, our approach essentially extends the synthetic controls ignorability assumption by imputing the counterfactuals absent treatment and conditioning on these quantities to obtain causal identification. In other words, while synthetic controls only conditions on the potential outcomes absent treatment prior to the implementation of treatment, our approach allows the analyst to condition on any subset of potential outcomes absent treatment up to the previous time-period, regardless of whether treatment occurred in the previous time-period. As a result, our ignorability assumption may be seen as an extension of an ignorability assumption underlying synthetic controls.
One key benefit of our approach relative to DiD or synthetic controls is the ability to estimate causal effects in the presence of dynamic policy landscapes where multiple policies change over time. In DiD or synthetic controls applications, the presence of other policy changes may be problematic. In DiD contexts, additional policy changes are essentially time-varying covariates. Because these policy changes likely differ in distribution between the treatment and control groups, these changes threaten the parallel-trends assumption and potentially confound DiD estimates (Zeldow and Hatfield, 2021). While identification may be recovered by conditioning on these policy changes as time-varying covariates, this requires a further assumption that the primary policy change of interest has no effect on other policy changes, which may not hold in practice. For synthetic controls applications, the presence of multiple changing policies does not necessarily threaten causal identification. However, the effect estimates will capture both a direct effect of the relevant policy on the outcome, and an indirect effect via the induced policy changes. As a result, the effect estimates capture effects of multiple policies. When the indirect effects are thought to be small relative to the direct effects, this may not be a problem; otherwise, these changes make the estimates harder to interpret. By contrast, our approach, when applying the extension to handle multiple policies, enables the identification of the direct effects of the policies, partialling out the effects of other policy changes. Moreover, unlike the DiD context, we require no assumption about the dependence of the policies to obtain identification.
In summary, our approach relies on an ignorability assumption that is similar to an ignorability assumption that may motivate synthetic controls; in fact, these assumptions may be identical in settings with only one treated period. The key difference is that our approach allows conditioning on post-treatment potential outcomes absent treatment, whereas synthetic controls only conditions on pre-treatment outcomes. Finally, in the presence of multiple policy changes, our approach can identify the direct effects a single policy change, whereas identification itself is threatened in a DiD context, and synthetic controls estimate will capture both direct and indirect effects, which may not be desirable.
@@ 4.2 The sequential g-formula
We now compare the identification result in Theorem 1 with the sequential g-formula (Gill and Robins, 2001). A more common approach in the epidemiology literature, the sequential g-formula relies on the so-called “sequential ignorability” assumption, which essentially states that in every time period, treatment is randomly assigned conditional on all prior outcomes, treatments, and covariates.  [^5] Sequential ignorability states that for all time-periods $t$, $Y_{it}(\bar{a}_{t})\perp A_{it}\mid\bar{Y}_{i,t-1},\bar{X}_{i,t-1},\bar{A}_{i,t-1}$., In combination with the corresponding positivity assumptions, sequential ignorability applied to our setting yields the identifying expression (the sequential g-formula),
@@
This expression is far more complex than the functional given in Theorem 1, which expresses our identified causal parameter in terms of a single averaged conditional expectation function. [^6] In practice, to estimate equation (5) structure is often placed on $\mathbb{E}[Y(\bar{a}_{t})]$; such models are known as marginal structural models (Robins, Hernan and Brumback, 2000). We obtain a simpler expression because our ignorability assumption conditions on the counterfactual outcomes absent treatment rather than the observed outcomes. To see this, consider a simple case with three time-periods, illustrated in Figure 1 below.
[…]
@@
However, equivalence between this expression and $\mathbb{E}[Y_{2}(\bar{a})]$ would require that $(A_{1},A_{2})\perp Y_{2}(a_{1}a_{2})\mid(Y_{0},Y_{1})$. Unfortunately, this will not hold under the causal structure depicted in Figure 1. The problem, as mentioned previously, is that $Y_{1}$ is a post-treatment outcome that acts as a collider between $A_{1}$ and $Y_{2}(a_{1}a_{2})$, since $Y_{1}$ lies on the causal pathway from $A_{1}$ to $Y_{2}$. Conditioning on $Y_{1}$ introduces a dependence between $Y_{2}(a_{1}a_{2})$ and $A_{1}$ that otherwise may not exist, violating the required ignorability assumption noted above.
By contrast, our ignorability assumption conditions on $(Y_{1}(00),Y_{0}(00))$, variables that do not lie on the causal pathway from $A_{1}$ to $Y_{2}$. Figure 2 depicts a causal diagram illustrating the assumed relationships between variables, where blue arrows indicate deterministic relationships (notice that the potential outcomes $Y_{t}(a_{1}a_{2})$, omitted from this diagram, are a deterministic function of $Y_{t}(00)$ by assumption). Figure 2 implies the conditional independence statement $(A_{1},A_{2})\perp Y_{2}(a_{1}a_{2})\mid Y_{1}(00),Y_{0}(00)$. By conditioning on the unobserved potential outcomes absent treatment rather than the observed outcomes, our approach does not introduce collider bias, and we are therefore able to achieve causal identification of an averaged conditional expectation function.
[…]

### v2

@@ 4.2 Additional time dependence
We may also wish to relax the no carryover effects (assumption 2) and ignorability assumptions (assumption 3), allowing for greater dependence between the treatments and outcomes over time. First consider no carryover effects: in many policy settings, treatment in earlier time periods likely affects outcomes at time $t$, and therefore, the effects of entire treatment sequences are of interest. To incorporate these effects we first redefine our causal estimand, now given by $\theta_{t}(\bar{a})=\mathbb{E}[Y_{it}(\bar{a})-Y_{it}(\bar{0})].$ This parameter represents the effect of a particular treatment sequence compared against no treatment at time $t$. While in principle this allows for effects of all prior treatments, in practice, it is often useful to make the assumption that these “carryover effects” are limited to the most recent $l$ time points.
@@
Let $\bar{a}$ and $\bar{a}^{\prime}$ be two treatment sequences such that $a_{t}=a_{t}^{\prime}$, $a_{t-1}=a_{t-1}^{\prime},\dots,a_{t-l}=a_{t-l}^{\prime}$, then we have $Y_{it}(\bar{a})=Y_{it}(\bar{a}^{\prime}),\ \forall\bar{a},\bar{a}^{\prime}.$
@@
Second, we relax Assumption 3 to condition on $k$ prior time periods’ counterfactuals as opposed to only the single time period’s counterfactual outcome used previously. This relaxation increases the plausibility of the assumption, particularly in settings where we worry about the dependence of the outcomes over time. However, in order to relax this assumption, we must strengthen the positivity assumption (Assumption 5) to include these additional time-periods. We may incorporate these relaxations into our estimation strategy by fitting the following version of the L-DAM:
$Y_{it}$ $=\alpha+\sum_{b=1}^{k}\left\{\beta_{b}\left(Y_{i,t-b}-\sum_{z=0}^{l}\theta_{z}A_{i,t-b-z}\right)\right\}+\sum_{z=0}^{l}\theta_{z}A_{i,t-z}+\epsilon_{it}.$
It is easy to see how this model incorporates these extensions. First, the $\theta_{z}$ parameters represent the lagged policy effects. Additionally, the model includes linear adjustments for estimates of $Y_{i,t-1}(0),\dots,Y_{i,t-k}(0)$ instead of estimates of only $Y_{i,t-1}(0)$.
@@ 4.3 Time-varying covariates
We can also relax our ignorability assumption to allow for both time-invariant and time-varying covariates. Assume, for example, that covariates $X_{it}$ are determined prior to $Y_{it}$. Then we could assume that ignorability holds conditional on covariates and $k$ lagged potential outcomes absent treatment $(X_{it}$, $Y_{it-k}(0),...,Y_{it-1}(0))$. However, including these covariates also requires a stronger positivity requirement to further condition on these quantities. One can then incorporate these same covariates into the L-DAM as:
$Y_{it}$ $=\alpha+\sum_{b=1}^{k}\left\{\beta_{b}\left(Y_{i,t-b}-\sum_{z=0}^{l}\theta_{z}A_{i,t-b-z}\right)\right\}+\sum_{z=0}^{l}\theta_{z}A_{i,t-z}+\Gamma^{\top}X_{it}+\epsilon_{it}.$
When the covariates $X_{it}$ are time-varying, causal identification also requires that these covariates are exogenous, or unaffected by treatment. That is, we must assume,
@@ 4.4 Multiple policies
A separate but important type of time-varying variable we may also observe are other policy indicators $M_{it}$. For example, state legislatures may consider multiple policies to tackle opioid-related outcomes, including both naloxone access laws and Medicaid prescription coverage policies. In this situation, we may wish to know the effect of one policy, holding the other policy fixed at some value. In the language of potential outcomes, we are interested in a so-called “controlled direct effect” (Nguyen, Schmid and Stuart, 2021).To be precise, assume that in each time period states may assign either policies $A$ and/or $M$, but we are primarily interested in the effects of policy $A$. One relevant causal quantity would then be $\psi_{t}(\bar{a},\bar{m})=\mathbb{E}[Y_{it}(\bar{a},\bar{m})-Y_{it}(\bar{0},\bar{m})]$, where potential outcomes are now indexed by both treatment values. Identifying this quantity in terms of the observed data distribution requires analogues to Assumptions (1)-(5). We formally describe these assumptions in Appendix A; however, intuitively we can view $(A,M)$ as a multi-valued treatment randomly assigned conditional on the prior potential outcomes absent any policy. [^5] We can allow that the randomization for $M$ is conditional on $A$ or vice versa. In much of the mediation literature it is assumed that $A$ is determined prior to $M$. We instead make no assumption on their causal ordering, at the price of a slightly stronger conditional independence assumption. Under these assumptions we can identify the causal estimand in terms of the observed data distribution as an average of a conditional expectation function, described in Appendix A. For continuous outcomes we can again parameterize a model for this function using an L-DAM:
$Y_{it}$ $=\alpha+\sum_{b=1}^{k}\left\{\beta_{b}\left(Y_{i,t-b}-\sum_{z=0}^{l}[\theta_{z}A_{i,t-b-z}+\zeta_{z}M_{i,t-b-z}]\right)\right\}+\sum_{z=0}^{l}[\theta_{z}A_{i,t-z}+\zeta_{z}M_{i,t-z}]+\epsilon_{it}$
@@ 4.5 Nonlinear extensions
Linear models may be inappropriate for modeling the conditional expectation function for certain outcomes. For example, in opioid studies, we often observe count data, where a log-link function may be more appropriate to model the conditional expectation function. We can continue to invoke Assumptions (1)-(5) above, but in place of assumption (4), we may instead assume constant multiplicative treatment effects:
@@
Analogous to the L-DAM, we can then parameterize the model using $\eta_{t}(\bar{a};\theta_{t})=\theta^{\top}\bar{a}$:
$\log(\mathbb{E}[Y_{it}\mid\bar{Y}_{i,t-1},\bar{A}_{it}])$ $=\alpha+\sum_{b=1}^{k}\left\{\beta_{b}\left(\log(Y_{i,t-b})-\sum_{z=0}^{l}\theta_{z}A_{i,t-1-z}\right)\right\}+\sum_{z=0}^{l}\theta_{z}A_{i,t-z}.$
@@
$\tilde{\theta}_{t}(\bar{a})=\frac{\mathbb{E}[Y_{it}(\bar{a})]}{\mathbb{E}[Y_{it}(\bar{0})]}.$
@@ 5 Connections to existing identification strategies
We briefly compare our identification strategy to other strategies commonly used in causal inference with panel data. First, we compare our approach to difference-in-differences and synthetic controls, commonly used in applied policy applications. Second, we compare it to the sequential g-formula, which identifies causal quantities under sequentially randomized trials and is a more commonly used approach in epidemiological studies.
@@ 5.1 Difference-in-differences and synthetic controls
Both difference-in-differences and synthetic control analyses are frequently motivated by assuming that the potential outcomes absent treatment follow a linear factor model,
$Y_{it}(0)=\bm{\lambda}_{t}^{\top}\bm{\mu}_{i}+\epsilon_{it},$
where $\bm{\lambda}_{t}$ is an $r$ by $1$ vector of so-called “factors” and $\bm{\mu}_{i}$ is an $r$ by $1$ vector of so-called “factor loadings”. Alternatively, one can think of $\bm{\lambda}_{t}$ as a time-varying coefficient vector applied to the unobserved covariate vector $\bm{\mu}_{i}$. Selection into treatment is then assumed also to be a function of the unobserved covariates $\bm{\mu}_{i}$. The two-way fixed effects model that often motivates difference-in-differences is a special case of this model where $r=2$, $\bm{\lambda}_{t}=[1,\lambda_{t}]$ and $\bm{\mu}_{i}=[\mu_{i},1]$. One may also extend this model to include observed covariates $X_{it}$ as done in (Abadie, Diamond and Hainmueller, 2010).
While we do not observe the unobserved confounder vector $\bm{\mu}_{i}$, we do observe the pre-treatment outcomes, which are a function of these confounders. Thus, the pre-treatment outcomes may serve as proxy variables for these unobserved confounders. Motivated by this insight, the synthetic controls method estimates the post-treatment counterfactual outcomes absent treatment for a treated unit by reweighting a set of control units to match the mean outcomes of the treated units in the pre-treatment period, and then using the weighted post-treatment outcomes of the control units for the estimate of the potential outcomes under control in the post-treatment period for the treatment group. Abadie, Diamond and Hainmueller (2010) show that under some conditions, as the number of pre-treatment time periods $T_{0}$ increases, the bias of the synthetic controls estimator approaches zero. Intuitively, the pre-treatment outcomes act as a proxy for the unobserved variables, and the more proxies we observe and adjust for, the better we adjust for the unobserved confounding.
We may similarly motivate our proposed method using the linear factor model, again taking the view that the potential outcomes absent treatment serve as proxies for unobserved confounders. However, in contrast to synthetic controls, our method uses both the pre- and post-treatment data to control for these variables. That is, we use both any observed pre-treatment outcomes and post-treatment estimates of the counterfactual outcomes absent treatment as proxies for these same variables. Assuming our models are well-specified, heuristically the bias of our estimator decreases with the number of lagged potential outcomes absent treatment – a function of $T$ – rather than the number of pre-treatment outcomes $T_{0}$. We prove this result in Proposition 5 in Appendix G, albeit under the assumption that the potential outcomes absent treatment are observed variables. One may therefore view our approach as a method to estimate causal effects under the linear factor model by imposing structure on the treatment effect to obtain post-treatment counterfactual estimates absent treatment.
Our proposed approach also has other benefits. In contrast to synthetic controls, our method more naturally accommodates multiple treated units and staggered treatment adoption. Additionally, in contrast to both synthetic controls and difference-in-differences, our method imputes both the potential outcomes absent treatment for the treated units – and the potential outcomes under treatment for the control units. This follows because our method imposes structure on the treatment effect that allows us to impute both counterfactuals. As a result, our method targets the average treatment effect (ATE), not simply the average treatment effect on the treated (ATT).
While the linear factor model is one way to motivate the synthetic controls approach, an alternative way is an ignorability assumption that conditions on pre-treatment outcomes (Ben-Michael, Feller and Rothstein, 2021; Abadie, Diamond and Hainmueller, 2010). This implies that treatment is randomized as a function of the pre-treatment outcomes at the time immediately before the first treatment period. Our ignorability assumption again simply generalizes this: we allow the potential outcomes absent treatment to enter the treatment assignment model, but make no restriction on whether these outcomes were actually observed – that is, whether they occurred pre- or post-treatment. Moreover, this also allows randomization to occur at every time period. In practice, this means that treatment assignment decisions are made with regard to what would have happened in the previous time periods had the intervention not taken place, whether or not the intervention was actually implemented.
@@ 5.2 The sequential g-formula
The sequential g-formula is a more common approach to estimating longitudinal treatment effects in the epidemiological literature, and relies on the so-called “sequential ignorability” assumption (Gill and Robins, 2001). This assumption states that in every time period, treatment is randomly assigned conditional on all prior outcomes, treatments, and covariates.  [^6] Sequential ignorability states that for all time-periods $t$, $Y_{it}(\bar{a}_{t})\perp A_{it}\mid\bar{Y}_{i,t-1},\bar{X}_{i,t-1},\bar{A}_{i,t-1}$. In combination with the corresponding positivity assumptions, sequential ignorability applied to our setting yields the identifying expression (the sequential g-formula),
@@
This expression is far more complex than our identification result, which expresses our identified causal parameter in terms of a single averaged conditional expectation function. [^7] In practice, to estimate equation (6) structure is often placed on $\mathbb{E}[Y(\bar{a}_{t})]$; such models are known as marginal structural models (Robins, Hernan and Brumback, 2000). We obtain a simpler expression because our ignorability assumption conditions on the counterfactual outcomes absent treatment rather than the observed outcomes. To see this, consider a simple case with three time-periods, illustrated in Figure 2 below.
$Y_{0}$$A_{1}$$Y_{1}$$A_{2}$$Y_{2}$U
@@
However, equivalence between this expression and $\mathbb{E}[Y_{2}(\bar{a})]$ would require that $(A_{1},A_{2})\perp Y_{2}(a_{1}a_{2})\mid(Y_{0},Y_{1})$. Unfortunately, this will not hold under the causal structure depicted in Figure 2. The problem, as mentioned previously, is that $Y_{1}$ acts as a collider between $A_{1}$ and $Y_{2}(a_{1}a_{2})$, since $Y_{1}$ lies on the causal pathway from $A_{1}$ to $Y_{2}$. Conditioning on $Y_{1}$ therefore introduces a dependence between $Y_{2}(a_{1}a_{2})$ and $A_{1}$, violating the required ignorability assumption.
By contrast, our ignorability assumption conditions on $(Y_{1}(00),Y_{0}(00))$, variables that do not lie on the causal pathway from $A_{1}$ to $Y_{2}$. Figure 3 depicts a causal diagram illustrating the assumed relationships between variables, where blue arrows indicate deterministic relationships (notice that the potential outcomes $Y_{t}(a_{1}a_{2})$, omitted from this diagram, are a deterministic function of $Y_{t}(00)$ by assumption). Figure 3 implies the conditional independence statement $(A_{1},A_{2})\perp Y_{2}(a_{1}a_{2})\mid Y_{1}(00),Y_{0}(00)$. By conditioning on the unobserved potential outcomes absent treatment rather than the observed outcomes, our approach avoids collider bias, allowing us to achieve causal identification in terms of an averaged conditional expectation function.
$Y_{0}$$A_{1}$$Y_{1}$$A_{2}$$Y_{2}$$Y_{0}(00)$$Y_{1}(00)$$Y_{2}(00)$$U$
@@
5.3 Model validation
Once a model has been specified, when sufficient pre-period data are available, we recommend conducting placebo tests similar to those used in the synthetic controls literature to choose or validate the specification (Abadie, Diamond and Hainmueller (2015)). Specifically, assume that we observe $t=1,...T_{0}$ time periods where no state is treated until time $T_{0}+1$. This leaves $T_{0}$ pre-treatment periods and $T_{P}=T-T_{0}$ post-treatment time periods. When $T_{0}\geq T_{P}$, we may take the observed distribution of treatment assignment in the post-treatment period and shift them backwards in time $T_{P}$ time-periods to generate pseudo-treatments in the pre-treatment period. When $T_{0}<T_{P}$, we could instead move either the first or final $T_{0}$ post-treatment time-periods back to the pre-treatment period. Finally, we may then run the model using this pre-period data and pseudo-treatment time periods to estimate the pseudo-treatment effects to test the models.
Specifically, when the pre-period estimates of $\theta$ are statistically indistinguishable from zero, this provides additional support to a given model. By contrast, a non-null effect may suggest a poorly specified model that may be revised, for example, by adding additional lagged outcomes or covariates to the specification. [^8] Notice, however, that using pre-treatment data to conduct model selection rather than model validation may result in overfitting. To conduct model selection, one may divide the pre-treatment data into two sets to first select then to validate the model (Abadie, Diamond and Hainmueller, 2015). The value of these placebo tests in this context, however, is slightly more limited than for synthetic controls or difference-in-differences applications. The goal of synthetic controls and DiD is to predict $Y_{it}(0)$ for the post-treatment treated units. However, our model must predict both $Y_{it}(0)$ for the post-treatment treated units and $Y_{it}(1)$ for the post-treatment control units. Heuristically, because we observe $Y_{it}(0)$ for the pre-treatment treated units, the pre-treatment data is arguably informative about the model of $Y_{it}(0)$ during the post-treatment period. However, this same logic does not hold for the post-treatment control units, whose values of $Y_{it}(1)$ are unobserved in the pre-treatment period. That is, because we do not observe $Y_{it}(1)$ for any units in the pre-treatment period, we also cannot model $Y_{it}(1)$ for any units, so that this placebo test does not inform any part of the model corresponding to the treatment effect. Despite this potential limitation, we believe these are still useful tests to implement and help users ensure their chosen statistical model is well-suited to their application.